OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 16?

OpenStudy (jhannybean):

equation of a circle: \((x-h)^2 +(y-k)^2 =r^2\)

OpenStudy (jhannybean):

\[(h,k) = (0,0)~,~ r^2 = 16 \implies r = \sqrt{16} = ~?\]